A chemist has 2.5 liters of a solution which is 70% acid. How much water should be added to obtain a 50% acid solution?
Answer by venugopalramana(3286)

chemist has 2.5 liters of a solution which is 70% acid. How much water should be added to obtain a 50% acid solution?
qty.of solution = 2.5 l
% acid in solution = 70
hence qty.of acid in the solution = 2.5*70/100 = 1.75 l
required percentage = 50
let the water to be added = x l
total qty.of solution = (2.5+x) l
qty of acid = same as before = 1.75 l
% acid in the new solution = 1.75*100/(2.5+x) = 50 as given
175= 50(2.5+x)
2.5+x = 175/50 = 3.5
x= 3.5 - 2.5 = 1 litre